# The Basic Reproduction Number for Petri Net Models: A Next-Generation Matrix Approach

### v2

@@ 4.6.1 Two Patch System
$\varphi{\tiny\begin{bmatrix}\frac{\left(\frac{\beta_{1}m_{{11}}p_{{11}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{12}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{1}}{\left(v_{1}+\mu_{1}\right)\left(\gamma_{1}+\delta_{1}+\mu_{1}\right)}&\frac{\left(\frac{\beta_{1}m_{{11}}p_{{12}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{22}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{2}}{\left(v_{2}+\mu_{2}\right)\left(\gamma_{2}+\delta_{2}+\mu_{2}\right)}&\frac{\frac{\beta_{1}m_{{11}}p_{{11}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{12}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{1}+\delta_{1}+\mu_{1}}&\frac{\frac{\beta_{1}m_{{11}}p_{{12}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{22}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{2}+\delta_{2}+\mu_{2}}\\ \frac{\left(\frac{\beta_{1}m_{{21}}p_{{11}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{12}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{1}}{\left(v_{1}+\mu_{1}\right)\left(\gamma_{1}+\delta_{1}+\mu_{1}\right)}&\frac{\left(\frac{\beta_{1}m_{{21}}p_{{21}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{22}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{2}}{\left(v_{2}+\mu_{2}\right)\left(\gamma_{2}+\delta_{2}+\mu_{2}\right)}&\frac{\frac{\beta_{1}m_{{21}}p_{{11}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{12}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{1}+\delta_{1}+\mu_{1}}&\frac{\frac{\beta_{1}m_{{21}}p_{{21}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{22}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{2}+\delta_{2}+\mu_{2}}\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}}$
@@
The dominant eigenvalue for this particular system is an incredibly long algebraic expression. However, with the $F$ and $V$ matrices matching that of the system outlined in the paper [23] and subsequently for two patches and two groups in Equations 32-39, the resulting $R_{0}$ coming from the dominant eigenvalue will be equal.
@@ 4.7 SIR Vector-borne Model
The outline of the original ODE system coming from Wedajo et al. describes the SIR Malaria model [24], but the structure of this model can be taken more broadly to represent a simple example of vector-borne disease modeling. The SIR vector-borne model demonstrates the adaptability of our method to different modes of transmission, extending its applicability from directly transmitted diseases to complex vector-host systems. The compartments $S_{h},I_{h},R_{h},S_{v},$ and $I_{v}$ represent susceptible humans, infected humans, recovered humans, susceptible vectors, and infected vectors. Note that in a vector-borne model, the vector’s infected compartments are considered general infected compartments for purposes of calculating $R_{0}$. The vector for Malaria in the original model is mosquitoes. The parameters $\Pi,\beta_{hv},\mu_{h},\delta,\alpha,\sigma,\Lambda,\beta_{vh},$ and $\mu_{v}$ represent birth/immigration of humans, the rate of infected vectors infecting humans, human natural death rate, the rate of infected humans immigrating, the rate of infected humans emigrating, the recovery rate, the birth/immigration rate of the vector, the rate of infected humans infecting vectors, and natural vector death rate, respectively.
$\frac{dS_{h}}{dt}$ $=$ $\Pi-\beta_{hv}S_{h}I_{v}-\mu_{h}S_{h}$ (40)
$\frac{dI_{h}}{dt}$ $=$ $\beta_{hv}S_{h}I_{v}+\delta I_{h}...$ (41)
$-(\alpha+\sigma+\mu_{h})I_{h}$
$\frac{dR_{h}}{dt}$ $=$ $\sigma I_{h}-\mu_{h}R_{h}$ (42)
$\frac{dS_{v}}{dt}$ $=$ $\Lambda-\beta_{vh}S_{v}I_{h}-\mu_{v}S_{v}$ (43)
$\frac{dI_{v}}{dt}$ $=$ $\beta_{vh}S_{v}I_{h}-\mu_{v}I_{v}$ (44)
Figure 8: SIR Vector-borne Variable Arc Weight PN Model
First, we obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the non-infected places token values, $(S_{h}^{*},S_{v}^{*},I_{h}^{*},I_{v}^{*},R_{h}^{*})=(\frac{\Pi}{\mu_{h}},\frac{\Lambda}{\mu_{v}},0,0,0)$. Then the matrix $\mathcal{F}$ by looking at the arc weights of arcs going into the infected compartments giving us
@@
From this, we find the rate of change with respect to $I_{h}$ in the first column and $I_{v}$ in the second column, and evaluate at the DFE to give
@@
$V$ $=$ $\begin{bmatrix}\alpha+\mu_{h}+\sigma-\delta&0\\ 0&\mu_{v}\end{bmatrix}$
@@
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}F\cdot V^{-1}\end{bmatrix}=\sqrt{(\frac{\beta_{hv}\Pi}{\mu_{h}\mu_{v}}\cdot\frac{\beta_{vh}\Lambda}{(\mu_{v})(\alpha+\mu_{h}+\sigma-\delta)})}$
Some of the parameters are renamed in the lay out of the ODE model in equations 40-44 versus the original [24]. With this parameter renaming, the resulting $R_{0}$ value equals the exact $R_{0}$ expression in the Wedajo paper.
@@ 5 Numerical Verification
This section presents the numerical verification of the algebraic expression for the basic reproduction number, $R_{0}$, derived from our Petri net $R_{0}$ methodology laid out above. While the analytical derivation of $R_{0}$ from the structural properties of the Petri net provides a concise mathematical representation of $R_{0}$ using the VAPN method, this section aims to provide robust, simulation-based evidence to corroborate these theoretical findings. In fact, the numerical results presented offer an independent validation of our $R_{0}$ expression.
@@
To achieve this validation, we conducted simulations of the Petri net model using GPenSim, a general-purpose Petri net simulator. GPenSim allows for close representation of ODE systems in Petri nets [7]. By analyzing the output of these simulations, specifically the progression of susceptible population and infectious population under various parameter sets, we can numerically estimate the $R_{0}$ exhibited by the simulated system. This analysis is independent of any algebraic $R_{0}$ methods that look at the parent system. The subsequent results will demonstrate the concordance between the $R_{0}$ value obtained through algebraic methods and the $R_{0}$ value inferred from the GPenSim simulation outcomes, thereby verifying the accuracy and reliability of our analytical expression. To show this we compare the $R_{0}$ values produced algebraically from the systems and through various methods of looking at the resulting simulation data. The comparison itself via relative root mean square error (RRMSE).
@@
The Petri net models and ODE models have been defined and run for a range of parameters. For models SIRS 4.1 and the Nonlinear model 4.5, error surfaces were found for the $R_{0}$ values depending on three of the parameters. The $R_{0}$ values from the respective ODE system and PN system are found directly from the resulting data using R-studio package called R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics library from Boelle and Obadia [25, 26]. Both the Attack Rate method [1] and the Maximum Likelihood method [27] within the library were utilized. The $R_{0}$ values for each system are then compared across the range of parameters to the equivalent algebraically, NGM and NGM-PN $R_{0}$ value.
5.1 SIRS
@@
The algebraic expression using the NGM and NGM-PN for the SIRS 4.1 systems is $R_{0}=\frac{\beta}{\gamma}$. The following Figure 9 compares the algebraic expression with that found from applying the attack rate method within the R library to the resulting simulation data across parameter values.
@@
Figure 9: Using the R library R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics from Boelle and Obadia [25, 26] with estimation method of $R_{0}$ of Attack rate. The error of the simulations remains below 1% RRMSE for all parameter values in the SIRS system and below 0.1% RRMSE for nearly all parameter values, including all biologically plausible values. Additionally, all $R_{0}$ values remain in the produced 95% confidence interval produced for the $R_{0}$ values given from the R library.
5.2 Nonlinear System
The algebraic expression of $R_{0}$ using the NGM and NGM-PN for the Nonlinear
refsec:NL systems is $R_{0}=R_{0}=\varrho(FV^{-1})=\frac{\sigma\beta}{(\sigma+\mu)(\gamma+\mu)}$. Since $R_{0}$ includes the parameters $\beta,\gamma,$ and $\sigma$ these will be used as the parameters where the $R_{0}$ values will be compared across a range for each.
@@ 5.2 Nonlinear System
Figure 10: Using the R library R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics from Boelle and Obadia [25, 26] with estimation method of $R_{0}$ of Attack rate. The error of the simulations remains below 1.2% RRMSE for all parameter values in the Nonlinear system and below 0.1% RRMSE for nearly all parameter values, including all biologically plausible values. Additionally, all $R_{0}$ values remain in the produced 95% confidence interval produced for the $R_{0}$ values given from the R library.
As shown in Figure 10, the Petri net implementation of the nonlinear model produces $R_{0}$ values that closely match the algebraic solution, with RRMSE consistently below 0.1%. This high level of agreement demonstrates the accuracy and numerical stability of the NGMPN method, even in the presence of nonlinear transmission and recovery terms. These results support the validity of the approach for more complex and mechanistically detailed epidemiological models.
@@ 6 Discussion and Conclusions
In this work, we devised a formal framework for finding the basic reproduction number $R_{0}$ within systems modeled by Petri nets. This framework draws upon well-established methodologies employed in the analysis of ODE models. The relevance of applying $R_{0}$ in the context of Petri nets is particularly pronounced, given the increasing utilization of Petri nets in epidemiological modeling to depict more intricate and dynamic scenarios of disease spread. These scenarios include multi-compartment models, spatially distributed populations, and frameworks incorporating social and nonlinear dynamics. Our approach builds on the next-generation matrix method, which is well understood within the realm of ODEs. Our results extend the definition of $R_{0}$ to Petri nets through the interpretation of token transitions and arc weights in a manner that aligns with underlying epidemiological processes.
@@
We demonstrate that the NGM methodology can be effectively employed in the context of Petri nets by first defining relevant matrices that represent the rate of infection $F$ and the rate of removal or transition out of compartments $V$. Subsequently, we compute the dominant eigenvalue of the product of $FV^{-1}$. This approach is consistent with established NGM techniques for ODE frameworks, resulting in a robust mechanism for quantifying the basic reproduction number when applied to Petri nets. The examples elucidated—including models such as SIRS, SEEIR, Basic COVID-19, Nonlinear, Patch, and SIR Vector-borne serve to illustrate the versatility of this methodology across a spectrum of disease dynamics characterized by various place to transition structures and arc weights.
@@
A primary challenge encountered in adapting the NGM framework to Petri nets pertains to effectively translating the compartmental structures and transition rates—typically articulated through arc weights and token movement—into a form compatible with the mathematical formulations employed in the NGM. Nevertheless, once the transition and removal rates are contextualized in terms of Petri net parameters using Diekmann’s approach [5], the ensuing process of calculating $R_{0}$ becomes straightforward and closely parallels that utilized for ODEs. This provides a powerful and generalizable method for scrutinizing disease dynamics in models based on Petri nets.
@@
Moreover, the framework afforded by Petri nets enables the integration of more complex features that are often challenging to model using traditional ODE approaches. These features include variable parameters, discrete events, and non-linear transitions [9]. Such flexibility renders Petri nets a compelling tool for accurately representing real-world epidemiological systems that exhibit complex interactions and dependencies, including co-infections, varying transmission rates, and spatial heterogeneity in disease dissemination.
@@
The proposed method, maintains practical limitations that also limit other methods of $R_{0}$ including those used for ODEs. These practical challenges persist in its application to extensive or highly intricate Petri nets. Specifically, the computational demands associated with calculating the dominant eigenvalue of large matrices can be substantial, particularly when addressing large-scale disease models that encapsulate numerous compartments. Again, these problems currently exist with all methods of algebraic $R_{0}$ formulation, but should be noted. Additionally, considerations pertaining to the discrete nature of token movement in Petri nets may necessitate careful deliberation when modeling diseases characterized by continuous dynamics.
@@
Future inquiries may be directed towards enhancing computational techniques to accommodate larger and more intricate systems, as well as exploring the application of this method to empirical epidemiological data. Overall, this work contributes to bridging the divide between traditional ODE-based modeling approaches and Petri net frameworks, thereby providing a more comprehensive toolkit for understanding and predicting disease propagation across a variety of contexts.
Funding
@@
Conflict of Interest
The authors declare no conflict of interest.

### v3

@@ 4.7 Patch System Overview
$R_{0}$ $=\varphi\begin{bmatrix}F\cdot V^{-1}\end{bmatrix}=\varphi{\small\begin{bmatrix}\frac{\left(\frac{\beta_{1}m_{{11}}p_{{11}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{12}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)\upsilon_{1}}{\left(\upsilon_{1}+\mu_{1}\right)\left(\gamma_{1}+\delta_{1}+\mu_{1}\right)}&\frac{\left(\frac{\beta_{1}m_{{11}}p_{{12}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{22}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)\upsilon_{2}}{\left(\upsilon_{2}+\mu_{2}\right)\left(\gamma_{2}+\delta_{2}+\mu_{2}\right)}&\frac{\frac{\beta_{1}m_{{11}}p_{{11}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{12}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{1}+\delta_{1}+\mu_{1}}&\frac{\frac{\beta_{1}m_{{11}}p_{{12}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{22}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{2}+\delta_{2}+\mu_{2}}\vskip 8.0pt\\ \frac{\left(\frac{\beta_{1}m_{{21}}p_{{11}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{12}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)\upsilon_{1}}{\left(\upsilon_{1}+\mu_{1}\right)\left(\gamma_{1}+\delta_{1}+\mu_{1}\right)}&\frac{\left(\frac{\beta_{1}m_{{21}}p_{{21}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{22}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)\upsilon_{2}}{\left(\upsilon_{2}+\mu_{2}\right)\left(\gamma_{2}+\delta_{2}+\mu_{2}\right)}&\frac{\frac{\beta_{1}m_{{21}}p_{{11}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{12}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{1}+\delta_{1}+\mu_{1}}&\frac{\frac{\beta_{1}m_{{21}}p_{{21}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{22}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{2}+\delta_{2}+\mu_{2}}\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}}$.
@@
The dominant eigenvalue for this particular system is an incredibly long algebraic expression. However, with the $F$ and $V$ matrices matching that of the system outlined in the paper Bichara and Iggidr (2018) and subsequently for two patches and two groups in Equations (35)–(42), the resulting $R_{0}$ coming from the dominant eigenvalue will be equal.
@@ 4.8 SIR Vector-Borne Model
The outline of the original ODE system, as described by Wedajo et al. Wedajo et al. (2018), is based on the SIR Malaria model. However, the structure of this model can be taken more broadly to represent a simple example of vector-borne disease modeling. The SIR vector-borne model demonstrates the adaptability of our method to different modes of transmission, extending its applicability from directly transmitted diseases to complex vector-host systems. The compartments $S_{h}$, $I_{h}$, $R_{h}$, $S_{v}$, and $I_{v}$ represent susceptible humans, infected humans, recovered humans, susceptible vectors, and infected vectors, respectively. Note that in a vector-borne model, the vector’s infected compartments are considered general infected compartments for purposes of calculating $R_{0}$. The vector for Malaria in the original model is mosquitoes. The parameters $\Pi,\beta_{hv},\mu_{h},\delta,\alpha,\sigma,\Lambda,\beta_{vh},$ and $\mu_{v}$ represent birth/immigration of humans, the rate of infected vectors infecting humans, human natural death rate, the rate of infected humans immigrating, the rate of infected humans emigrating, the recovery rate, the birth/immigration rate of the vector, the rate of infected humans infecting vectors, and natural vector death rate, respectively. The SIR Vector-borne system is described by the following system of equations:
$\frac{dS_{h}}{dt}$ $=\Pi-\beta_{hv}S_{h}I_{v}-\mu_{h}S_{h},$ (43)
$\frac{dI_{h}}{dt}$ $=\beta_{hv}S_{h}I_{v}+\delta I_{h}\ldots$
$-(\alpha+\sigma+\mu_{h})I_{h},$ (44)
$\frac{dR_{h}}{dt}$ $=\sigma I_{h}-\mu_{h}R_{h},$ (45)
$\frac{dS_{v}}{dt}$ $=\Lambda-\beta_{vh}S_{v}I_{h}-\mu_{v}S_{v},$ (46)
$\frac{dI_{v}}{dt}$ $=\beta_{vh}S_{v}I_{h}-\mu_{v}I_{v}.$ (47)
From Figure 9 we determine that there are two infected places ($I_{H}$, $I_{v}$) per patch. Then, $F$ and $V$ matrices will be two by two.
First, we obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the non-infected places token values, $(S_{h}^{*},S_{v}^{*},I_{h}^{*},I_{v}^{*},R_{h}^{*})=(\frac{\Pi}{\mu_{h}},\frac{\Lambda}{\mu_{v}},0,0,0)$. Then the matrix $\mathcal{F}$ is obtained by looking at the arc weights of arcs going into the infected compartments, giving us
@@
Then we find the rate of change with respect to $I_{h}$ in the first column and $I_{v}$ in the second column, and plugging in the DFE to give
@@
We then find the rate of change with respect to the infected place $I$, and evaluate at the DFE. From this, we find the inverse to give
$V$ $=\begin{bmatrix}\alpha+\mu_{h}+\sigma-\delta&0\\ 0&\mu_{v}\end{bmatrix}$,
@@
Then the $F$ and $V$ matrices, allow us to determine $R_{0}$. Thus,
$R_{0}$ $=\varphi\begin{bmatrix}F\cdot V^{-1}\end{bmatrix}=\sqrt{(\frac{\beta_{hv}\Pi}{\mu_{h}\mu_{v}}\cdot\frac{\beta_{vh}\Lambda}{(\mu_{v})(\alpha+\mu_{h}+\sigma-\delta)})}.$
Some of the parameters are renamed in the layout of the ODE model in Equations (43)–(47) versus the original Wedajo et al. (2018). With this parameter renaming, the resulting $R_{0}$ value equals the exact $R_{0}$ expression in the Wedajo paper.
Figure 9: A variable arc weight SIR Vector-borne PN mapped from Equations (43)–(47). The arc weights used in forming matrix $F_{i}(x)$ are outlined in red, and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
@@ 5 Numerical Verification
This section presents the numerical verification of the algebraic $R_{0}$ expressions derived from our framework. This validation is performed by comparing the analytical $R_{0}$, found using the NGMPN method, to a numerically estimated $R_{0}$ derived from simulation output. We specifically focus on simulation output from the deterministic Variable Arc Weight Petri Net (VAPN) formalism. The VAPN’s ability to be numerically convergent with standard ODEs has been systematically studied, establishing it as a valid, deterministic PN implementation for this type of comparison Reckell et al. (2024). Therefore, our numerical verification scheme is a direct, deterministic comparison, for which RRMSE is the most appropriate measure of error. While our NGMPN framework also applies to stochastic SPN models, numerical validation with SPN models would involve stochastic analysis, which would be suitable for future work. The numerical results presented offer an independent validation of our $R_{0}$ expression.
@@
To achieve this validation, we conducted simulations of the Petri Net model using GPenSim, a general-purpose Petri Net simulator. GPenSim allows for close representation of ODE systems in Petri Nets Reckell et al. (2024). By implementing the VAPN and analyzing the output of these simulations, specifically the progression of the susceptible and infectious populations under various parameter sets, we can numerically estimate the $R_{0}$ exhibited by the simulated system. This analysis is independent of any algebraic $R_{0}$ methods that look at the parent system. The subsequent results will demonstrate the concordance between the $R_{0}$ value obtained through algebraic methods and the $R_{0}$ value inferred from the GPenSim simulation outcomes, thereby verifying the accuracy and reliability of our analytical expression. To illustrate this, we compare the $R_{0}$ values obtained algebraically from the systems with those derived through various methods of analyzing the resulting simulation data. The comparison is made via the relative root mean square error (RRMSE).
@@
The Petri Net models and ODE models have been defined and run for a range of parameters. For models SIRS Section 4.1 and the Nonlinear model Section 4.6, error surfaces were found for the $R_{0}$ values depending on three of the parameters. The $R_{0}$ values from the VAPN system are found directly from the resulting simulation data using the R package “R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics” Version:1.3-1 library from Boelle and Obadia Obadia et al. (2012); Boelle and Obadia . The Maximum Likelihood (ML) White et al. (2009) method within the package was utilized to estimate $R_{0}$ from the simulation time series. This estimated $R_{0}$ is then compared to the analytically derived $R_{0}$ value from our NGMPN method. The error is quantified using RRMSE. We find strong agreement in numerical terms because the Petri Net $R_{0}$ calculations achieve RRMSE values consistently below 1–2% for biologically plausible parameter sets, as shown in Figures 10 and 11. While the R packages Obadia et al. (2012); Boelle and Obadia can produce confidence intervals (CIs) for stochastic data, our primary validation here is a deterministic-to-deterministic comparison, making RRMSE the most direct and rigorous metric for goodness of fit.
@@
The algebraic expression using the NGM and NGMPN for the SIRS Section 4.1 systems is $R_{0}=\frac{\beta N}{\gamma}$. The following Figure 10 compares the algebraic expression with that found from applying the attack rate method within the R library to the resulting simulation data across parameter values. For the Nonlinear system in Section 4.6, the algebraic expression of $R_{0}$ using the NGM and NGMPN is $R_{0}=\varrho(FV^{-1})=\frac{\sigma\beta}{(\sigma+\mu)(\gamma+\mu)}$. Since $R_{0}$ includes the parameters $\beta$, $\gamma$, and $\sigma$, these will be used as the parameters for comparing $R_{0}$ values across a range for each.
@@
Figure 10: Using the R library “R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics” from Boelle and Obadia Obadia et al. (2012); Boelle and Obadia with estimation method of $R_{0}$ of Attack rate. The error of the simulations remains below 1% RRMSE for all parameter values in the SIRS system and below 0.1% RRMSE for most parameter sets and below 1% RRMSE for all parameter sets, including all biologically plausible values.
@@
Figure 11: Using the R library R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics from Boelle and Obadia Obadia et al. (2012); Boelle and Obadia with estimation method of $R_{0}$ of ML. The error of the simulations remains below 1.2% RRMSE for all parameter values in the Nonlinear system and below 0.1% RRMSE for nearly all parameter values, including all biologically plausible values.
As shown in Figure 11, the Petri Net implementation of the nonlinear model produces $R_{0}$ values that closely match the algebraic solution, with RRMSE consistently below 0.1% for most parameter sets and below 1% for all parameter sets. This high level of agreement demonstrates the accuracy and numerical stability of the NGMPN method, even in the presence of nonlinear transmission and recovery terms. These results support the validity of the approach for more complex and mechanistically detailed epidemiological models.
@@ 6 Discussion and Conclusions
In this work, we devised a formal framework for finding the basic reproduction number $R_{0}$ within systems modeled by Petri Nets. This framework draws upon well-established methodologies employed in the analysis of ODE models. The relevance of applying $R_{0}$ in the context of Petri Nets is particularly pronounced, given the increasing utilization of Petri Nets in epidemiological modeling to depict more intricate and dynamic scenarios of disease spread. These scenarios encompass multi-compartment models, spatially distributed populations, and frameworks that incorporate social and nonlinear dynamics. Our approach builds on the next-generation matrix method, which is well understood within the realm of ODEs. Our results extend the definition of $R_{0}$ to Petri Nets by interpreting token transitions and arc weights in a manner that aligns with the underlying epidemiological processes. As seen in Section 3.2, by focusing on a formal, algorithmic identification of arcs corresponding to new infections, $F$, contrasted with other transitional changes, $V$, our framework offers a direct path to calculating $R_{0}$. This formal clarity enables the ’by-inspection’ application and color-coding in our figures, which provides a direct visual representation of our method’s partition logic.
The primary contributions of our paper are threefold. First, we provide a clear and formalism–agnostic operational framework for constructing the NGM. It applies directly to Petri Nets regardless of their structure, and we demonstrate its effectiveness on both traditional SPNs and the more flexible VAPNs. Second, we apply this method to a wide range of complex, contemporary models, including multi-strain COVID-19 and multi-patch systems. Lastly, we offer the first, to our knowledge, numerical validation of an NGMPN method, showing that our analytical results align precisely with $R_{0}$ values estimated from simulation data. This validation is a crucial step in ensuring that these theoretical tools are reliable for practical application in epidemiological modeling.
@@
In parallel to the work performed by Segovia (2025), we demonstrate that the NGM methodology can be effectively employed in the context of Petri Nets by first defining relevant matrices that represent the rate of infection, $F$, and the rate of removal or transition out of compartments, $V$. Subsequently, we compute the dominant eigenvalue of the product of $FV^{-1}$. This approach is consistent with established NGM techniques for ODE frameworks, resulting in a robust mechanism for quantifying the basic reproduction number when applied to Petri Nets. The examples we laid out, including models such as SIRS, SVEIR, SEEIR, Basic COVID-19, Nonlinear, Patch, and SIR Vector-borne, serve to illustrate the versatility of this methodology across a spectrum of disease dynamics characterized by various place to transition structures and arc weights.
@@
A primary challenge encountered in adapting the NGM framework to Petri Nets pertains to effectively translating the compartmental structures and transition rates—typically articulated through arc weights and token movement—into a form compatible with the mathematical formulations employed in the NGM. Nevertheless, once the transition and removal rates are contextualized in terms of Petri Net parameters using Diekmann’s approach Diekmann et al. (2010), the ensuing process of calculating $R_{0}$ becomes straightforward and closely parallels that utilized for ODEs. This provides a robust and generalizable method for scrutinizing disease dynamics in models based on Petri Nets.
@@
Moreover, the framework afforded by Petri Nets enables the integration of more complex features that are often challenging to model using traditional ODE approaches. These features include variable parameters, discrete events, and nonlinear transitions Peleg et al. (2005). The SVEIR model in Section 4.4 demonstrates how this framework naturally extends beyond the static $R_{0}$ to calculate the time-dependent effective reproduction number $R_{e}$, incorporating features like vaccination rates Van den Driessche and Watmough (2002); Cori et al. (2013). Such flexibility renders Petri Nets a compelling tool for accurately representing real-world epidemiological systems that exhibit complex interactions and dependencies, including co-infections, varying transmission rates, and spatial heterogeneity in disease dissemination.
@@
The proposed method maintains practical limitations that also limit other methods of $R_{0}$, including those used for ODEs. These practical challenges persist in their application to extensive or highly intricate Petri Nets. Specifically, the computational demands associated with calculating the dominant eigenvalue of large matrices can be substantial, particularly when addressing large-scale disease models that encapsulate numerous compartments. Again, these problems currently exist with all methods of algebraic $R_{0}$ formulation, but they should be noted. Additionally, considerations pertaining to the discrete nature of token movement in Petri Nets may necessitate careful deliberation when modeling diseases characterized by continuous dynamics. For SPNs in particular, the deterministic $R_{0}=1$ threshold has known explanatory limitations, as it does not fully capture the probability of stochastic fadeout where an outbreak dies out by chance even if $R_{0}>1$, especially with small initial numbers of infected individuals. This “gap” between the deterministic threshold and stochastic behavior can be bridged by considering the system’s scale. As demonstrated numerically in Reckell et al. (2024), SPN models show numerical convergence to their corresponding deterministic ODE and VAPN counterparts as the total population size is rescaled to larger values. While not a theoretical proof, this numerical convergence indicates that the deterministic $R_{0}$ derived from our framework is at minimum a practical approximation for the limit for the stochastic process. Therefore, while our $R_{0}$ value does not predict stochastic fadeout in small populations, it remains the crucial and accurate metric for predicting the mean, large-scale outbreak potential of the stochastic system. A more systematic analysis of the probabilistic behavior around this $R_{0}=1$ threshold for different population scales is an important area for future investigation.
@@
Future work may be directed towards enhancing computational techniques to accommodate larger and more intricate systems, as well as exploring the application of this method to empirical epidemiological data. Overall, this work contributes to bridging the divide between traditional ODE-based modeling approaches and Petri Net frameworks, thereby providing a more comprehensive toolkit for understanding and predicting disease propagation across various contexts.
[…]
Conceptualization: T.R., B.S. and P.J.; Methodology: T.R.; Software, T.R.; Validation: T.R.; Formal Analysis: T.R.; Investigation: T.R.; Resources: T.R.; Data Curation: T.R.; Writing—original draft preparation: T.R.; Writing—review and editing, T.R., B.S. and P.J.; Visualization: T.R.; Supervision, B.S. and P.J.; Project Administration: B.S. and P.J.; Funding Acquisition: B.S. and P.J. All authors have read and agreed to the published version of the manuscript.
[…]
@@
[…]
Not applicable.
[…]
Not applicable.
[…]
Data and code used to produce data used in this paper is available in the publicly accessible repository GitHub at https://github.com/trevorreckell/NGMPN (accessed on 10 October 2025).
Acknowledgements.
[…]